# Generalized Face Liveness Detection via De-spoofing Face Generator

### v1

@@ 4.3 Leave-One-Out Experiments
Following the Leave-One-Out protocol with the same settings in [19], we each time use three of the four datasets (MSU-MFSD, CASIA-FASD, Idiap Replay-Attack, and OULU-NPU) for training, while the remaining one dataset is used for testing. We compare our method with the state-of-the-art (SOTA) methods and the experimental results are summarized in Table 1.
It can be observed that our method outperforms all the existing SOTA methods. We attribute this to the utilization of the anomalous cue that guides the learning of the FAS feature. The anomalous cue is learned by our DFG from a real face dataset whose data diversity far exceeds the existing FAS datasets, which allows the cue to provide robust hints about the anomaly of the input in different scenarios. Consequently, the proposed AG-Net catches more generalized attack representations that are effective in wider domains.
4.4 Experiments with Limited Source Domains
Table 2: Results with limited source domains
CASIA OULU Mean
HTER AUC HTER AUC HTER AUC
MADDG [17] 41.02 64.33 39.35 65.10 40.19 64.72
SSDG-M [19] 31.89 71.29 36.01 66.88 33.95 69.09
D2AM [55] 32.65 72.04 27.70 75.36 30.18 73.70
SSAN-M [23] 30.00 76.20 29.44 76.62 29.72 76.41
Resnet18 [54] 20.56 85.42 21.39 87.43 20.97 86.43
SSDG-R [19] 20.56 85.42 21.39 87.43 20.97 86.43
SSAN-R [23] 27.33 79.52 28.75 77.78 28.04 78.65
SA-FAS [35] 30.11 73.87 28.25 80.02 29.18 76.94
AG-FAS (Ours) 7.33 96.96 14.36 92.98 10.85 94.97
[…]
@@ 4.4 Experiments with Limited Source Domains
We conducted cross-domain FAS experiments in a more challenging setting with limited-source domains. We use the MSU-MFSD and the Idiap Replay-Attack datasets for training and the rest two datasets for testing. The results are shown in Table 2. Despite the limited availability of training data, our method outperforms all existing SOTA methods. This outcome demonstrates that the anomalous cue can still guide the FAS model to learn a more generalized feature even with limited source training data.
4.5 Ablation Study
Table 3: Comparisons between different feature extraction methods
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
only residual 20.00 85.23 33.44 70.82 13.43 93.96 23.89 84.67 22.69 83.67
dual-branch 7.14 96.86 7.22 98.05 9.29 96.31 13.32 94.02 9.24 96.31
AG-Net (Ours) 5.71 97.76 5.89 98.14 5.86 98.36 9.41 96.36 6.72 97.65
[…]
We conduct experiments to study the effectiveness of the obtained anomalous cue and the structure of the AG-Net in our method, and the results are summarized in Table 3. We first only feed the original image of the FAS dataset into the model without introducing any anomalous cues in the setting “only image”. In other words, the setting “only image” actually represents our implementation of the SSDG [19] method with the ViT-base [51] as the backbone. This setting is taken as the baseline.
The setting “only residual” replaces the image inputs in the setting “only image” with the anomalous cues obtained by the DFG, and the model structures of these two settings are the same. As observed in the results, the “only residual” setting that solely uses anomalous cues yields inferior performance compared to the “only image” setting that solely uses the original images. The reason is that though the anomalous cue can to some extent indicate where anomaly is present in the input, it cannot fully capture all the attack representations in the original image. The result demonstrates that the anomalous cue should be used together with the original image to assist FAS feature learning.
We then try to adopt the dual-branch architecture to combine the information of the anomalous cue and the original image, denoting the “dual-branch” setting. The dual-branch network uses two ViT-base models to extract the features of the anomalous cue and the original image, respectively, and concatenates the features of them before feeding it to the softmax classification layer. Although the “dual-branch” setting generally shows a slight improvement compared to the baseline, the result is not satisfactory since the FAS feature extractor essentially does not benefit from the anomalous cue during the simple feature combination.
In contrast, our proposed AG-Net leverages the anomalous cue throughout the entire FAS feature extraction process via cross-attention modules, which keeps providing hints for the feature extractor to search for more generalized attack representations. By effectively utilizing the anomalous cue, our method obtains the best performance on the Leave-One-Out experiments.
4.6 Different Approaches for Obtaining the Anomalous Cue
Table 4: Ablation study for different anomalous cue obtaining approaches
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
baseline 6.91 97.06 8.00 97.37 7.71 97.96 16.11 92.43 9.68 96.20
MemAE [41] 5.95 97.99 5.33 98.43 9.50 96.44 10.54 96.25 7.83 97.28
Deep SVDD [38] 5.95 97.42 6.67 98.06 8.36 97.49 12.95 94.71 8.48 96.92
AnoDDPM [45] 8.57 96.88 6.56 98.22 11.64 95.81 9.46 96.60 9.06 96.88
DFG (Ours) 5.71 97.76 5.89 98.14 5.86 98.36 9.41 96.36 6.72 97.65
[…]
We conduct experiments to compare our DFG model with other commonly used autoencoder-based anomaly detection methods, determining which one extracts more effective anomalous cues. Specifically, we train MemAE [41], Deep SVDD [38], and a DDPM-based anomaly detection method [45] on the same real face dataset. We then use these methods to reconstruct “real” version images on the FAS dataset and calculate the corresponding anomalous cue. For MemAE and Deep SVDD, we can directly get the reconstructed “real” version image by feeding the input to the network. For DDPM, which generates random images from Gaussian noise, we follow the approach presented in AnoDDPM [44] by adding noise to the input image for only 200 steps and then using DDPM to denoise it, thus obtaining the “real” version image corresponding to the input.
Subsequently, we use the anomalous cues obtained by the methods above as the input of our AG-Net, respectively, and conduct the FAS experiments. The training setting is consistent with the Leave-One-Out protocol. We still take the baseline in Table 3 without any cue guidance as the baseline in this experiment. The results in Table 4 show that the proposed DFG-based anomalous cue guides the FAS model to obtain the best overall performance. Compared with other anomalous cue obtaining methods, the DFG method obtains a cue that performs more stable on all the settings, which means our method is more reliable when faced with different scenarios.
The visualization of the reconstruction results obtained by the methods above on the FAS dataset is shown in Figure 3.
Figure 3: Visualization of the reconstructed images obtained by different methods and their corresponding residual with the input images.
It can be observed that MemAE and Deep SVDD have relatively poor generating capabilities as the reconstructed images contain obvious generated patterns of distortion even with real faces as input. Specifically, the images generated by MemAE exhibit noticeable blocky artifacts, while the residuals obtained from Deep SVDD display evident horizontal and vertical stripes. Although the DDPM-based method exhibits good reconstruction results on real faces, the reconstructed spoof faces with presentation attacks also closely resemble the input, making the anomalous cues of the real and spoof faces hard to distinguish.
In comparison to the aforementioned methods, the proposed DFG achieves promising results. As for the real images, the reconstructed faces are almost identical to the input. As for the spoof images, the reconstructed faces show a reduction in the artifacts present in the inputs, such as printing patterns, printing color differences, paper edges, reflections, and brightness anomalies, which makes the generated images exhibit differences from the inputs. This allows the anomalous cue obtained from DFG to effectively assist the FAS model in distinguishing between real and spoof faces.
4.7 Study the Impact of Introducing Additional Real Faces
@@
Table 5: Comparison with existing DG-based methods that use additional real faces (*: using additional real faces during training)
@@
SSDG-R* [19] 8.57 96.27 18.00 89.58 14.50 94.43 13.06 94.53 13.53 93.71
SSAN-R* [23] 9.76 95.81 22.67 85.48 22.14 86.40 15.14 92.80 17.43 90.12
SA-FAS* [35] 13.10 92.37 25.33 83.51 22.86 78.45 21.11 86.17 20.60 85.13
Ours* 5.71 97.76 5.89 98.14 5.86 98.36 9.41 96.36 6.72 97.65
[…]
@@
Since the proposed AG-FAS leverages additional real faces to help improve the FAS performance, we want to investigate whether existing DG-based methods can also utilize these data to improve their performance. Specifically, we include the 200,000 additional real faces into the training dataset of the DG-based methods. To ensure balanced data sampling, we make the model sample real data from the original FAS dataset and the additional real faces with the same probability. The results of the experiments are summarized in Table 5. It can be observed that all the DG-based methods experience a performance decline when additional real faces are added to the training dataset. This might be attributed to the gap in the FAS feature distribution between the additional real faces and the test FAS dataset. It is difficult to improve the model generalization by simply introducing additional real faces to existing DG-based methods. On the contrary, our AG-FAS learns an anomalous cue from the real faces, which is a more general pattern across different scenarios and thus helps improve the cross-dataset performance of the model, demonstrating the superiority of tackling the model generalization problem.
4.8 Generalization to Other Datasets
Table 6: A Larger-scale FAS cross-domain benchmark
CelebA-Spoof HKBU-MARs WFFD Rose-Youtu SiW Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
SSDG-R [19] 23.73 85.36 23.53 84.22 37.74 65.29 12.22 94.54 4.97 98.68 20.44 85.62
SSAN-R [23] 24.81 83.31 22.12 85.68 40.24 63.49 14.15 92.80 8.85 97.01 22.03 84.46
SA-FAS [35] 29.30 77.54 31.63 74.86 41.34 60.67 16.42 90.33 6.79 98.01 25.10 80.28
AG-FAS (Ours) 16.90 90.36 18.37 90.56 33.53 70.00 10.45 96.17 2.38 99.68 16.33 89.35
[…]
@@ 4.8 Generalization to Other Datasets
To further validate the effectiveness of AG-FAS, we conduct experiments on a broader range of commonly used datasets in the field of FAS to test the generalization ability of our approach. Specifically, we train different DG-based models on MSU-MFSD, CASIA-FASD, Idiap Replay-Attack, and OULU-NPU and then evaluate the performance of these models on other commonly used FAS datasets, i.e., CelebA-Spoof, HKBU-MARs, WFFD, Rose-Youtu, and SiW, respectively. The experimental results are presented in Table 6. As can be seen, our method achieves the best performance on all the testing FAS datasets, further demonstrating that anomalous cue guides the model to catch more generalized attack representations in the face images.
@@
It is worth noting that among the FAS datasets used for testing, the HKBU and WFFD datasets include 3D mask and wax figure presentation attacks, respectively, which are not contained in the four datasets used for training. The experimental results on these two datasets prove that our method also performs well in the challenging FAS task with unknown presentation attacks.
5 Conclusion

### v2

@@ V-C Leave-One-Out Experiments
Following the Leave-One-Out protocol with the same settings in [21], we each time use three of the four datasets (MSU-MFSD, CASIA-FASD, Idiap Replay-Attack, and OULU-NPU) for training, while the remaining one dataset is used for testing. We compare our method with the state-of-the-art (SOTA) methods and the experimental results are summarized in Table I. It can be observed that our method outperforms all the existing SOTA methods. We attribute this to the utilization of the anomalous cue that guides the learning of the FAS feature.
Note that DDM also employs the idea of reconstructing real faces, using residuals for liveness detection. However, because DDM is not designed for the domain generalization task, its performance is relatively poor. Therefore, we first introduce the SSDG strategy into DDM so it can better perform the DG tasks. Then, to ensure a fair comparison, we replace the backbone of DDM with the same one as ours. The improved DDM method is denoted as “DDM*” in the table.
The improved DDM exhibits good performance on this DG protocol, but it is still inferior to our method. This is because we utilize the face identity as the guidance for the DFG model, which allows us to add noise for more steps in the forward process according to the discussion in our theoretical analysis. As a result, a larger time step leads to better elimination of spoof information in the input, thus making anomalous cues more distinguishable. Besides, compared to DDM which learns the feature from two separate branches for the input image and the residual, our OA-Net enables the network to better focus on the off-real region according to the anomalous cue via the cross-attention modules, thereby obtaining features more relevant to the fake patterns.
TABLE II: Results with limited source domains
M&I to C M&I to O
HTER AUC HTER AUC
MADDG [19] 41.02 64.33 39.35 65.10
SSDG-M [21] 31.89 71.29 36.01 66.88
D2AM [65] 32.65 72.04 27.70 75.36
SSAN-M [25] 30.00 76.20 29.44 76.62
SSDG-R [21] 20.56 85.42 21.39 87.43
SSAN-R [25] 27.33 79.52 28.75 77.78
SA-FAS [35] 30.11 73.87 28.25 80.02
AG-FAS (Ours) 7.33 96.99 15.97 91.10
V-D Experiments with Limited Source Domains
@@
We conducted cross-domain FAS experiments in a more challenging setting with limited-source domains. We use MSU-MFSD and Idiap Replay-Attack datasets for training and the rest two datasets for testing. The results are shown in Table II. Despite the limited availability of training data, our method outperforms all existing SOTA methods. This outcome demonstrates that the anomalous cue can still guide the FAS model to learn a more generalized feature even with limited source training data.
V-E Study DG-based Methods with Additional Data
@@
TABLE III: Comparison with existing DG-based methods that use additional real faces (*: using additional real faces during training)
@@
SSDG-R* [21] 10.00 95.26 17.33 89.07 4.93 99.28 11.96 94.94 11.06 94.64
SSAN-R* [25] 11.43 94.80 21.33 87.54 12.86 94.08 19.65 89.65 16.32 91.52
SA-FAS* [35] 25.48 82.05 24.44 84.42 22.86 84.76 25.04 82.96 24.45 83.55
AG-FAS*(Ours) 5.71 98.03 5.44 98.55 6.714 98.23 9.43 96.62 6.82 97.86
Since the proposed AG-FAS leverages additional real faces to help improve the FAS performance, we want to investigate whether existing DG-based methods can also benefit from these data. Specifically, we include the 130,000 additional real faces in the training dataset of the DG-based methods. To ensure balanced data sampling, in addition to maintaining an overall 1:1 ratio between real and fake faces during sampling, we also resample the real faces within each minibatch. This ensures a 1:1 ratio between the real faces from the original FAS dataset and the additional real faces. The results of the experiments are summarized in Table III. It can be observed that the performance of the DG-based methods does not show a significant improvement; instead, there is even a noticeable degradation in some methods.
@@
This might be attributed to the gap in the FAS feature distribution between the additional real faces and the test FAS dataset. It is difficult to improve the model generalization by simply introducing additional real faces to existing DG-based methods. On the contrary, our AG-FAS learns an anomalous cue from the real faces, which is a more general pattern across different scenarios and thus helps improve the cross-dataset performance of the model, demonstrating the superiority of tackling the model generalization problem.
TABLE IV: Ablation study on different anomalous cue obtaining approaches (†: mask the background of the anomalous cue)
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
baseline 6.91 97.06 8.00 97.37 7.71 97.96 16.11 92.43 9.68 96.20
MemAE [49] 4.29 98.48 6.11 98.12 8.36 97.54 11.94 95.35 7.67 97.37
Deep SVDD [46] 5.71 97.97 6.67 97.82 9.29 97.57 11.81 95.14 8.37 97.13
AnoDDPM [55] 7.14 97.29 7.33 97.72 6.79 97.79 12.08 95.51 8.34 97.08
DFG†(Ours) 5.95 97.39 4.00 98.70 7.07 97.63 10.28 96.17 6.83 97.47
DFG (Ours) 5.71 98.03 5.44 98.55 6.71 98.23 9.43 96.62 6.82 97.86
V-F Different Approaches for Obtaining the Anomalous Cue
We conduct experiments to compare our DFG model with other commonly used autoencoder-based anomaly detection methods, determining which one extracts more effective anomalous cues. Specifically, we train MemAE [49], Deep SVDD [46], and a DDPM-based anomaly detection method [55] on the same real face dataset. We then use these methods to reconstruct “real” version images on the FAS dataset and calculate the corresponding anomalous cue. For MemAE and Deep SVDD, we directly get the reconstructed “real” version image by feeding the input to the network. For DDPM, the image-based diffusion model, we follow the approach presented in AnoDDPM [52] by adding noise to the input image for only 200 steps and then using DDPM to denoise it, thus obtaining the “real” version image corresponding to the input.
Subsequently, we use the anomalous cues obtained by the methods above as the input of our OA-Net, respectively, and conduct the FAS experiments. The training setting is consistent with the Leave-One-Out protocol. We take the ViT-base model without any cue guidance as the baseline in this experiment. The results in Table IV show that all methods that introduce the anomalous cues generally exhibit improvement compared to the baseline and the proposed DFG-based anomalous cue guides the FAS model to obtain the best overall performance. Compared with other methods, the DFG method obtains a cue that performs more stable on all the settings, which means our method is more reliable when faced with different scenarios.
We find that the reconstructed background in the anomalous cues may differ from the input, even for real images, which is due to DFG’s lack of additional information about the input background. To investigate the role of background in the anomalous cues, we conduct comparative experiments by masking the background regions of the anomalous cues, with the results denoted as “DFG†” in Table IV. The background mask is obtained by processing the RGB input using the BiSeNet [67] network. The results show that removing the background regions does not significantly affect the overall experimental outcomes. We attribute this to the fact that the anomalous cues are passed through a trainable cue encoder in the OA-Net, enabling the model to adaptively focus on the information relevant to the presentation attacks while diminishing its emphasis on the noise caused by the background reconstruction.
V-G Visualization of the Anomalous Cue
The visualization of the reconstruction results obtained by the methods above on the FAS dataset is shown in LABEL:fig:reconstruction. To better highlight the differences in the anomalous cues obtained by each method in the facial region, we present the residual images after masking the background regions. It can be observed that MemAE and Deep SVDD have relatively poor generating capabilities as the reconstructed images contain obvious generated patterns of noise. This noise present in the generation process can affect the effectiveness of the anomalous cue. Although the DDPM-based method exhibits relatively good reconstruction results, the reconstructed spoof faces with presentation attacks also closely resemble the input, making the anomalous cues of the real and spoof faces hard to distinguish.
In comparison to the aforementioned methods, the proposed DFG achieves promising results. As for the real images, the reconstructed faces are almost identical to the input. As for the spoof images, the reconstructed faces show a reduction in the artifacts present in the inputs, such as printing patterns and brightness anomalies, which makes the generated images exhibit differences from the inputs. This allows the anomalous cue obtained from DFG to effectively assist the FAS model in distinguishing between real and spoof faces.
V-H Ablation Study
V-H1 Comparisons between Different Structures of OA-Net
TABLE V: Comparisons between different structures of OA-Net
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
only residual 17.14 88.44 26.44 82.42 21.57 89.00 17.36 90.92 20.63 87.69
dual-branch 7.14 97.19 6.67 98.16 11.36 95.39 13.07 94.34 9.56 96.27
OA-Net (Ours) 5.71 98.03 5.44 98.55 6.71 98.23 9.43 96.62 6.82 97.86
We conduct experiments to study the effectiveness of the obtained anomalous cue and the structure of the OA-Net in our method, and the results are summarized in Table V. We still take the ViT-base model without any cue guidance as the baseline in this experiment, denoting the “only image” setting.
The setting “only residual” replaces the image inputs in the setting “only image” with the anomalous cues obtained by the DFG. As observed in the results, the “only residual” setting that solely uses anomalous cues yields inferior performance compared to the “only image” setting that solely uses the original images. The reason is that though the anomalous cue can to some extent indicate where anomaly is present in the input, it cannot fully capture all the attack representations in the original image. The result demonstrates that the anomalous cue should be used together with the original image to assist FAS feature learning.
We then try to adopt the dual-branch architecture to combine the information of the anomalous cue and the original image, denoting the “dual-branch” setting. The dual-branch network uses two ViT-base models to extract the features of the anomalous cue and the original image, respectively, and concatenates the features before feeding them to the softmax classification layer. The result of the “dual-branch” setting is not satisfactory since the FAS feature extractor essentially does not benefit from the anomalous cue via a simple concatenation of the features.
In contrast, our proposed OA-Net leverages the anomalous cue throughout the entire FAS feature extraction process via cross-attention modules, which keeps providing hints that guide the feature extractor to learn from the off-real regions. By effectively utilizing the anomalous cue, our method obtains the best performance on the Leave-One-Out experiments.
TABLE VI: Performance with the anomalous cue generated under different time steps
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
$\hat{t}=1000$ 5.71 97.84 7.22 97.62 11.36 95.33 8.19 97.50 8.12 97.07
$\hat{t}=800$ 5.71 98.03 5.44 98.55 6.71 98.23 9.43 96.62 6.82 97.86
$\hat{t}=600$ 5.48 98.50 6.67 98.06 8.50 97.13 8.21 97.23 7.21 97.73
$\hat{t}=400$ 4.05 98.41 7.44 97.58 7.64 97.55 10.57 95.88 7.43 97.35
$\hat{t}=200$ 7.38 97.35 6.00 98.40 8.71 97.30 8.59 96.90 7.67 97.49
V-H2 Effectiveness of the Identity Feature
We investigate the influence of the condition identity feature on the training of DFG. We train a DFG model without the identity feature and compare it with a DFG model trained under standard settings. From LABEL:fig:id_feature, it can be seen that, without the identity feature as the condition, the DFG tends to exhibit changes in identity while generating the “real” face. In contrast, the DFG model trained with identity guidance keeps the consistency of the identity in the generated face across different diffusion steps.
Additionally, we find that using only the identity feature ($\hat{t}=1000$) to guide the generation of the “real” face still preserves the pose and expression of the people in the generated images. This suggests that the employed Arcface encoder extracts additional information, such as the pose, beyond the identity feature.
V-H3 Impact of the Diffusion Steps
We conduct ablation studies on the diffusion steps while generating the anomalous cue, and the results are shown in Table VI. We find that the model’s performance tends to decline when the diffusion steps are either too high or too low. When the diffusion steps are set as $\hat{t}=1000$, the diffusion model reconstructs the image from a latent code in random noise, which introduces more noise into the anomalous cue. However, when the number of the diffusion steps is low, the diffusion model tends to generate images identical to the input, thereby diminishing the attack information in the anomalous cue.
TABLE VII: Performance with the DFG trained on different data quantity
O&C&I to M O&M&I to C O&C&M to I I&C&M to O Mean
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
25% 6.91 97.71 5.44 98.42 8.79 96.46 11.96 95.03 8.27 96.91
50% 5.95 97.93 6.67 97.78 7.14 97.93 9.39 96.93 7.29 97.64
100% 5.71 98.03 5.44 98.55 6.71 98.23 9.43 96.62 6.82 97.86
V-H4 Impact of the Quantity of the Additional Real Faces
We conduct ablation experiments on the quantity of additional real faces to study its impact on the proposed method. Specifically, we randomly select a certain proportion of the real faces for the training of our DFG and evaluate the performance. The results are shown in Table VII. It can be observed that with an increase in the data used for DFG, the overall performance of the model improves. This suggests that when we provide more data of real faces during training, the proposed DFG can actually gain better knowledge of what a real face should be like and thus generate more effective anomalous cues. The result above indicates there is potential for future improvement in the performance of AG-FAS if more real faces can be obtained.
V-I Generalization to Other Datasets
TABLE VIII: A Larger-scale FAS cross-domain benchmark
HKBU-MARs WFFD Rose-Youtu SiW WMCA
HTER AUC HTER AUC HTER AUC HTER AUC HTER AUC
SSDG-R [21] 23.53 84.22 37.74 65.29 12.22 94.54 4.97 98.68 14.61 91.98
SSAN-R [25] 22.12 85.68 40.24 63.49 14.15 92.80 8.85 97.01 16.79 90.40
SA-FAS [35] 31.63 74.86 41.34 60.67 16.42 90.33 6.79 98.01 18.91 88.08
AG-FAS (Ours) 17.98 90.43 35.44 67.23 11.77 95.13 3.18 99.50 9.95 96.84
@@ V-I Generalization to Other Datasets
To further validate the effectiveness of AG-FAS, we conduct experiments on a broader range of commonly used datasets in the field of FAS to test the generalization ability of our approach. Specifically, we train different DG-based models on MSU-MFSD, CASIA-FASD, Idiap Replay-Attack, and OULU-NPU and then evaluate the performance of these models on other commonly used FAS datasets, i.e., HKBU-MARs, WFFD, Rose-Youtu, SiW, and WMCA, respectively. The experimental results are presented in Table VIII. As can be seen, our method achieves the best performance on all the testing FAS datasets, further demonstrating that anomalous cue guides the model to catch more generalized attack representations in the face images.
@@
It is worth noting that among the FAS datasets used for testing, the HKBU, WFFD, and WMCA datasets include 3D mask and wax figure presentation attacks, respectively, which are not contained in the four datasets used for training. The experimental results on these two datasets prove that our method also performs well in the challenging FAS task with unknown presentation attacks.
VI Conclusion